OpenStudy (anonymous):

(5x^3-4)/(5x^4+1) differentiate

OpenStudy (anonymous):

i got ((5x^4) (15x^2) - (5x^3)(20x^3)) / (5x^4+1)^2 but it is suppose to be in the form (i am suppose to fill in the letters): ((-25x^A) + (Bx^3) + (Cx^D)) / (5x^4+1)^F

OpenStudy (anonymous):

u=5x^3-4 u'=15x^2 v=5x^4+1 v'=20x^3 dy/dx = vu'-uv' / v^2 dy/dx = (5x^4+1)(15x^2)-(20x^3)(5x^3-4) / (5x^4+1)^2 dy/dx = (75x^6+15x^2)-(100x^6-80x^4) / (5x^4+1)^2 dy/dx = (-25x^6+80x^4+15x^2) / (5x^4+1)^2

OpenStudy (anonymous):

how did you go from the second to last step to the last step?

OpenStudy (anonymous):

second step to last step ? this is quotient rule 1st we will take numerator as u 2nd we will take denominator as v 3rd differentiate both to get u' n v' 4th use its formula, which is dy/dx = vu'-uv' / v^2 5th expand your numerator then, by referring to your question, you're required to compare A, B and C with your answer, well, you got the value, what left now is simply take it, and compare the value of A, B and C with those who have same term :)

OpenStudy (anonymous):

o i just didnt expand properly thx

OpenStudy (anonymous):

can you help me with another one please :):)